OpenStudy (anonymous):

what is the surface area of a ice cream cone r= 5 s= 36

OpenStudy (lgbasallote):

hint: \[SA = \pi rs + \pi r^2\] where: r = radius s = slant height does that help?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Right

Directrix (directrix):

Assuming that the ice cream cone is not sealed, the problem is to find the lateral area of the right circular cone. There is no visible base for which to find area. Lateral Area = pi * r * s where r is the radius and s is the slant height. LA = pi * 5 * 36 = 180 square units.